# Vortical waves in a quantum fluid with vector, axial and helical charges. II. Dissipative effects

### v1

@@ 6 Summary and Conclusions
For the environment of the quark-gluon plasma produced in the current experiments in relativistic heavy-ion collisions, the estimations made in Section 5 do not favour the observation of the Helical Vortical Waves. It appears that the angular velocity of the plasma is too small, so the wave propagates too slowly (about 1% of the speed of light), implying that during the lifetime of the plasma fireball, the wave advances over a phenomenologically irrelevant distance of a small fraction of a fermi. However, it appears that even this estimation is too optimistic because of the damping effects. The helical relaxation time sets an upper bound of the wavelength of the propagating wave because when the wavelengths are too large, wave propagation gets restricted as a result of the non-conservation of the helical charge. However, even this wavelength appears to be too small as the kinetic dissipation effects – that increase with the decrease of the wavelength – appear to be too strong so that they damp the wave efficiently. Therefore, we conclude that helical vortical waves cannot propagate in the realistic environment of quark-gluon plasmas.
It is worth mentioning that we do not consider the effect of finite electric conductivity that promotes a diffusive electromagnetic backreaction of an oscillating electric charge density generated by the waves. This electromagnetic effect can be relevant for all hydrodynamic excitations that involve oscillations of the vector chemical potential. An analysis of the electromagnetic backreaction has been performed for the Chiral Magnetic Wave in the background magnetic field in Ref. Shovkovy:2018tks and for the longitudinal sound wave, a circularly polarized vortical wave and diffusive modes, among others, in a vortical background in Ref. Rybalka:2018uzh . It appears that these waves are overdamped in quark-gluon plasma, which aligns well with our conclusions made for the helical and axial hydrodynamic excitations. Moreover, a finite electric conductivity will further limit the propagation of helical and axial waves as they also contain a vector component. Therefore, our theoretical analysis gives an upper bound on the dissolution time of these modes. On the other hand, the electromagnetic effects are irrelevant to the charge-neutral Axial Vortical excitations, which will be discussed in a separate publication.
Acknowledgments. This work is supported by the European Union - NextGenerationEU through grant No. 760079/23.05.2023, funded by the Romanian ministry of research, innovation and digitalization through Romania’s National Recovery and Resilience Plan, call no. PNRR-III-C9-2022-I8.
@@
In order to see that the two approaches are inequivalent, let us consider the variation of $Q_{H}$ with respect to small perturbations in $T$ and $\mu_{V/A/H}$. For simplicity, we restrict ourselves to the large temperature expression for $Q_{H}$. From Eq. (20) with $\ell=H$ (see also Eq.(3.5b) in Ref. HVWp1 ), we find that
$Q_{H}=\frac{\mu_{H}T^{2}}{3}+\dfrac{4T\ln 2}{\pi^{2}}\mu_{A}\mu_{V}+\dfrac{\mu_{H}(3\mu_{V}^{2}+3\mu_{A}^{2}+\mu_{H}^{2})}{3\pi^{2}}+O(T^{-1})$ (159)
@@
$\delta Q_{H}{\biggl{|}}_{\mu_{A}=\mu_{H}=0}=\left(\frac{T^{2}}{3}+\frac{\mu_{V}^{2}}{\pi^{2}}\right)\delta\mu_{H}+\frac{4\ln 2}{\pi^{2}}\mu_{V}T\delta\mu_{A}+O(T^{-1}).$ (160)
Therefore, in the prescription for relaxation given by $\nabla_{\mu}J^{\mu}_{H}=-Q_{H}/\tau_{H}$ there is an extra term coming from the crossed susceptibility $\chi_{HA}=\partial Q_{H}/\partial\mu_{A}$ which is responsible for a relaxation towards a state with $Q_{H}=0$, which does not necessarily enforce $\mu_{H}=0$. On the contrary, the prescription used in the main text $\nabla_{\mu}J^{\mu}_{H}=-T^{2}\mu_{H}/3\tau_{H}$ does lead to a relaxation towards $\mu_{H}=0$. We proceed now to discuss the consequences of introducing dissipation on the basis of charge.
@@
[…]
The matrix $\mathbb{M}$ is now slightly different from that of section 3. In particular:
$\dfrac{k}{T^{2}}\mathbb{M}=\omega\mathbb{M}_{\omega}-\kappa_{\Omega}\mathbb{M}_{\Omega}+\dfrac{2i\sigma_{A}^{\omega}}{\tau_{H}T^{2}}\mathbb{I}_{H}+\dfrac{2i\sigma_{H}^{\omega}}{\tau_{H}T^{2}}(\delta_{\ell H}\delta_{\ell^{\prime}A})\,.$ (162)
In this case, the matrix $\mathbb{M}$ in Eq. (45) changes to
$\frac{1}{T^{2}}\mathbb{M}_{\ell\ell^{\prime}}=\omega\mathbb{M}_{\omega}-\kappa_{\Omega}\mathbb{M}_{\Omega}+\dfrac{2i\sigma_{A}^{\omega}}{\tau_{H}T^{2}}\mathbb{I}_{H}+\dfrac{2i\sigma_{H}^{\omega}}{\tau_{H}T^{2}}(\delta_{\ell H}\delta_{\ell^{\prime}A}).$ (163)
Its determinant can be evaluated as follows:
${\rm det}\left(\frac{1}{T^{2}}\mathbb{M}\right)={\rm det}\left(\omega\mathbb{M}_{\omega}-\kappa_{\Omega}\mathbb{M}_{\Omega}+\dfrac{2i\sigma_{A}^{\omega}}{\tau_{H}T^{2}}\mathbb{I}_{H}\right)-\frac{2i\sigma_{H}^{\omega}}{\tau_{H}T^{2}}\left|\begin{array}[]{cc}\frac{2\omega}{T^{2}}\left(\sigma_{A}^{\omega}-\frac{T^{2}}{3}\Delta H\right)&-\frac{\kappa_{\Omega}}{H}B\\ -\kappa_{\Omega}A&\frac{2\omega}{T^{2}}\sigma_{H}^{\omega}\end{array}\right|,$ (164)
@@
Accordingly, the master equation ${\rm det}(\mathbb{M})=0$ gets modified to
$\frac{2\omega}{T^{2}}\left\{\left(\frac{2\omega}{T^{2}}\right)^{2}\left(\sigma^{\omega}_{A}-\frac{T^{2}}{3}\Delta H\right)[(\sigma^{\omega}_{A})^{2}-(\sigma^{\omega}_{H})^{2}]-\frac{\kappa_{\Omega}^{2}}{H}[(A^{2}+B^{2})\sigma_{A}^{\omega}-2AB\sigma^{\omega}_{H}]\right\}\\ +\frac{i}{3\tau_{H}}\left[\left(\frac{2\omega}{T^{2}}\right)^{2}\dfrac{6}{T^{2}}\left(\sigma_{A}^{\omega}-\frac{T^{2}}{3}\Delta H\right)[(\sigma_{A}^{\omega})^{2}-(\sigma_{H}^{\omega})^{2}]-\frac{\kappa_{\Omega}^{2}A}{H}\dfrac{6}{T^{2}}(A\sigma_{A}^{\omega}-B\sigma_{H}^{\omega})\right]=0\,.$ (165)
We can solve the previous equation perturbatively in $\kappa_{\Omega}$. Let us expand $\omega$ in powers of $\kappa_{\Omega}$:
@@
$\left\{\left(\frac{2\omega_{0}}{T^{2}}\right)^{2}\left(\sigma^{\omega}_{A}-\frac{T^{2}}{3}\Delta H\right)[(\sigma^{\omega}_{A})^{2}-(\sigma^{\omega}_{H})^{2}]\right\}\left(\frac{2\omega_{0}}{T^{2}}+\dfrac{2i}{\tau_{H}T^{2}}\right)=0\,.$ (167)
There is a non-trivial and purely dissipating solution $\omega_{0}=-i/\tau_{H}$ and a double trivial root $\omega_{0}=0$. We now compute the corrections to the trivial root from the next to the leading order for the master equation:
$\dfrac{2i}{\tau_{H}T^{2}}\left[\left(\frac{2\tau_{H}\omega_{1}}{T^{2}}\right)^{2}\left(\sigma_{A}^{\omega}-\frac{T^{2}}{3}\Delta H\right)[(\sigma_{A}^{\omega})^{2}-(\sigma_{H}^{\omega})^{2}]-\frac{A}{H}(A\sigma_{A}^{\omega}-B\sigma_{H}^{\omega})\right]\kappa_{\Omega}^{2}=0$ (168)
@@
$\omega_{1}^{\pm}=\pm\dfrac{T^{2}}{2\tau_{H}}\sqrt{\dfrac{A(A\sigma_{A}^{\omega}-B\sigma_{H}^{\omega})}{H\left(\sigma_{A}^{\omega}-\frac{T^{2}}{3}\Delta H\right)[(\sigma_{A}^{\omega})^{2}-(\sigma_{H}^{\omega})^{2}]}}\,.$ (169)
The quantity $A(A\sigma_{A}^{\omega}-B\sigma_{H}^{\omega})$ is negative semi-definite, whereas the remaining factor inside the square root is positive semi-definite. Specifically, from the definitions, we find $\sigma_{A}^{\omega}>\sigma_{H}^{\omega}$ and
$H=\frac{15\alpha_{V}^{4}+30\pi^{2}\alpha_{V}^{2}+7\pi^{4}}{15\alpha_{V}^{2}+7\pi^{2}}>0,\quad\frac{1}{T^{2}}\left(\sigma_{A}^{\omega}-\frac{T^{2}}{3}\Delta H\right)=\frac{15\alpha_{V}^{4}+6\alpha_{V}^{2}\pi^{2}+7\pi^{4}}{90\alpha_{V}^{2}+42\pi^{2}}>0,$ (170)
@@
$\dfrac{1}{T^{2}}A(A\sigma_{A}^{\omega}-B\sigma_{H}^{\omega})\simeq-0.000432\alpha_{V}^{2}<0$ (171)
@@
Figure 13: Explicit demonstration that the left hand side of Eq. (171) is negative definite and therefore there is always one unstable mode at small wavenumber.
@@ Appendix B Helicity relaxation time at finite chemical potential
In this section of the appendix, we extend the calculation of the helicity relaxation time $\tau_{H}$ presented Ref. Ambrus:2019khr to the case of an unpolarized, charged plasma, i.e. at finite vector chemical potential $\mu_{V}=\alpha_{V}T$. We start from Eqs. (101), (102) and (104) of Ref. Ambrus:2019khr :
@@
[…]
×$\times[f^{\sigma}_{{\mathbf{p}}^{\prime},-\lambda}f^{-\sigma}_{{\mathbf{k}}^{\prime},\lambda}\tilde{f}^{\sigma}_{{\mathbf{p}},\lambda}\tilde{f}^{-\sigma}_{{\mathbf{k}},-\lambda}-f^{\sigma}_{{\mathbf{p}},\lambda}f^{-\sigma}_{{\mathbf{k}},-\lambda}\tilde{f}^{\sigma}_{{\mathbf{p}}^{\prime},-\lambda}\tilde{f}^{-\sigma}_{{\mathbf{k}}^{\prime},\lambda}],$ (172)
with $s=E^{2}_{\rm cm}=(p+k)^{2}=(p^{\prime}+k^{\prime})^{2}$ being the center-of-mass energy and $\theta_{\rm cm}$ the angle between ${\mathbf{p}}^{\prime}$ and ${\mathbf{p}}$, measured in the center of mass frame, satisfying
@@
where $f^{\sigma}_{0{\mathbf{p}}}=[e^{\beta E_{\mathbf{p}}-\sigma\alpha_{V}}+1]^{-1}$ is the equilibrium distribution for a charged, unpolarized fluid, thus extending Eq. (105) of Ref. Ambrus:2019khr for the neutral fluid to the case of a charged fluid. Note that in Eq. (107) of Ref. Ambrus:2019khr , Eq. (172) was replaced by $dQ_{H}/dt=-Q_{H}/\tau_{H}$, which is valid under the assumption of a neutral background with a small helicity imbalance, i.e. $\mu_{V}=\mu_{A}=0$ and $|\mu_{H}|\ll T$, when $Q_{H}\simeq\mu_{H}T^{2}/3$. In this work, we consider the case when $\mu_{V}$ is arbitrary, such that instead Eq. (172) becomes
@@
where we used the property $\tilde{f}^{\sigma}_{0{\mathbf{p}}}=1-f^{\sigma}_{0{\mathbf{p}}}$. Writing now as in Eq. (159),
@@
Figure 14: Ratio $\tau_{H}(0)/\tau_{H}(\alpha_{V})$ between the helicity relaxation time in a neutral plasma and at dimensionless vector chemical potential $\alpha_{V}=\mu_{V}/T$, in an unpolarized plasma.
The integral with respect to $\varphi_{{\mathbf{p}}^{\prime}}$ can be performed using the explicit expression for $1-\cos\theta_{\rm cm}$,
$1-\cos\theta_{\rm cm}=\frac{4p\cdot p^{\prime}}{E_{\rm cm}^{2}}.$ (182)
After boosting $p^{\prime}$ to the center of mass frame, $p^{\mu}$ is replaced by $p^{\mu}_{\rm cm}=(E_{\rm cm}/2,{\mathbf{p}}_{\rm cm})$, with
@@
Splitting ${\mathbf{p}}_{\rm cm}={\mathbf{p}}_{||}+{\mathbf{p}}_{\perp}$, where ${\mathbf{p}}_{||}$ is oriented along ${\mathbf{k}}+{\mathbf{p}}$ and ${\mathbf{p}}_{\perp}$ is perpendicular to this vector, we have
$p_{||}=\frac{{\mathbf{p}}_{\rm cm}\cdot({\mathbf{p}}+{\mathbf{k}})}{|{\mathbf{p}}+{\mathbf{k}}|}=\frac{E_{\rm cm}(p^{0}-k^{0})}{2|{\mathbf{p}}+{\mathbf{k}}|},$ (184)
@@
$\tau_{H}=0.196\times\frac{6\pi^{3}\beta}{g\alpha_{\rm QCD}^{2}}\simeq\left(\frac{250\ {\rm MeV}}{k_{B}T}\right)\left(\frac{1}{\alpha_{\rm QCD}}\right)^{2}\left(\frac{2}{N_{f}}\right)\times 4.804\ {\rm fm}/c.$ (187)
@@
$\tau_{H}^{-1}(\alpha_{V}\gg 1)=c_{0}\frac{e^{-\alpha_{V}}}{\alpha_{V}}\,$ (188)
where the best fit was found to be given by $c_{0}=118.6\,$.

### v2

@@ 6 Summary and Conclusions
For the environment of the quark-gluon plasma produced in the current experiments in relativistic heavy-ion collisions, the estimations made in Section 5 do not favour the observation of the Helical Vortical Waves. It appears that the angular velocity of the plasma is too small, so the wave propagates too slowly (about 1% of the speed of light), implying that during the lifetime of the plasma fireball, the wave advances over a phenomenologically irrelevant distance of a small fraction of a fermi. However, it appears that even this estimation is too optimistic because of the damping effects. The helical relaxation time sets an upper threshold of the wavelength of the propagating wave because when the wavelengths are too large, wave propagation is prohibited as a result of the non-conservation of the helical charge. However, even this threshold wavelength appears to be too small as the kinetic dissipation effects – that increase quadratically with the decrease of the wavelength – appear to be too strong so that they damp the wave at small wavelengths. Therefore, we conclude that helical vortical waves cannot propagate in the realistic environment of quark-gluon plasmas.
It is worth mentioning that we do not consider the effect of finite electric conductivity that promotes a diffusive electromagnetic backreaction of the oscillating electric charge density generated by the waves. This electromagnetic effect can be relevant for all hydrodynamic excitations that involve oscillations of the vector chemical potential. An analysis of the electromagnetic backreaction has been performed for the Chiral Magnetic Wave in the background magnetic field in Ref. Shovkovy:2018tks and for the longitudinal sound wave, a circularly polarized vortical wave and diffusive modes, among others, in a vortical background in Ref. Rybalka:2018uzh . It appears that these waves are overdamped in quark-gluon plasma, which aligns well with our conclusions made for the helical and axial hydrodynamic excitations. Moreover, a finite electric conductivity will further limit the propagation of helical and axial waves as they also contain a vector component. Therefore, our theoretical analysis gives an upper bound on the dissolution time of these modes.
Acknowledgments. We thank P. Aasha for fruitful discussions. We also thank an anonymous referee for helping us improve the connections to existing literature. This work was funded by the EU’s NextGenerationEU instrument through the National Recovery and Resilience Plan of Romania - Pillar III-C9-I8, managed by the Ministry of Research, Innovation and Digitization, within the project entitled “Facets of Rotating Quark-Gluon Plasma” (FORQ), contract no. 760079/23.05.2023 code CF 103/15.11.2022.
@@
In order to see that the two approaches are inequivalent, let us consider the variation of $Q_{H}$ with respect to small perturbations in $T$ and $\mu_{V/A/H}$. For simplicity, we restrict ourselves to the large temperature expression for $Q_{H}$. From Eq. (19) with $\ell=H$ (see also Eq. (64b) in Ref. Morales-Tejera:2024uzg ), we find that
$Q_{H}=\frac{\mu_{H}T^{2}}{3}+\dfrac{4T\ln 2}{\pi^{2}}\mu_{A}\mu_{V}\\ +\dfrac{\mu_{H}(3\mu_{V}^{2}+3\mu_{A}^{2}+\mu_{H}^{2})}{3\pi^{2}}+O(T^{-1})$ (159)
@@
$\delta Q_{H}{\biggl{\rvert}}_{\mu_{A}=\mu_{H}=0}=\left(\frac{T^{2}}{3}+\frac{\mu_{V}^{2}}{\pi^{2}}\right)\delta\mu_{H}\\ +\frac{4T\ln 2}{\pi^{2}}\mu_{V}\delta\mu_{A}+O(T^{-1}).$ (160)
Therefore, in the prescription for relaxation given by $\partial_{\mu}J^{\mu}_{H}=-Q_{H}/\tau_{H}$ there is an extra term coming from the crossed susceptibility $\chi_{HA}=\partial Q_{H}/\partial\mu_{A}$ which is responsible for a relaxation towards a state with $Q_{H}=0$, which does not necessarily enforce $\mu_{H}=0$. On the contrary, the prescription used in the main text $\partial_{\mu}J^{\mu}_{H}=-T^{2}\mu_{H}/3\tau_{H}$ does lead to a relaxation towards $\mu_{H}=0$. We proceed now to discuss the consequences of introducing dissipation on the basis of charge.
@@
$\partial_{\mu}J^{\mu}_{V}=0\,,\hskip 8.5359pt\partial_{\mu}J^{\mu}_{A}=0\,,\hskip 8.5359pt\partial_{\mu}J^{\mu}_{H}=-\dfrac{Q_{H}}{\tau_{H}}\,.$ (161)
The matrix $\mathbb{M}$ is now slightly different from that in Eq. (15a). In particular, taking the $\tau_{A}\to\infty$ limit, we obtain:
$\frac{1}{T^{2}}\mathbb{M}_{\ell\ell^{\prime}}=\omega\mathbb{M}^{\omega}_{\ell\ell^{\prime}}-\kappa_{\Omega}\mathbb{M}^{\Omega}_{\ell\ell^{\prime}}+\dfrac{2i\sigma_{A}^{\omega}}{\tau_{H}T^{2}}\delta_{\ell H}\delta_{\ell^{\prime}H}\\ +\dfrac{2i\sigma_{H}^{\omega}}{\tau_{H}T^{2}}(\delta_{\ell H}\delta_{\ell^{\prime}A}),$ (162)
where the last term is due to the cross-susceptibility $\chi_{HA}$. The determinant of $\mathbb{M}$ can be evaluated as follows:
${\rm det}\left(\frac{1}{T^{2}}\mathbb{M}\right)={\rm det}\left(\omega\mathbb{M}_{\omega}-\kappa_{\Omega}\mathbb{M}_{\Omega}+\dfrac{2i\sigma_{A}^{\omega}}{\tau_{H}T^{2}}\mathbb{I}_{H}\right)\\ -\frac{2i\sigma_{H}^{\omega}}{\tau_{H}T^{2}}\left\lvert\begin{array}[]{cc}\frac{2\omega}{T^{2}}\left(\sigma_{A}^{\omega}-\frac{T^{2}}{3}\Delta H\right)&-\frac{\kappa_{\Omega}}{H}B\\ -\kappa_{\Omega}A&\frac{2\omega}{T^{2}}\sigma_{H}^{\omega}\end{array}\right\rvert,$ (163)
@@
Accordingly, the master equation ${\rm det}(\mathbb{M})=0$, shown in Eq. (33), gets modified to
$\frac{2\omega}{T^{2}}\left\{\left(\frac{2\omega}{T^{2}}\right)^{2}\left(\sigma^{\omega}_{A}-\frac{T^{2}}{3}\Delta H\right)[(\sigma^{\omega}_{A})^{2}-(\sigma^{\omega}_{H})^{2}]\right.\\ \left.-\frac{\kappa_{\Omega}^{2}}{H}[(A^{2}+B^{2})\sigma_{A}^{\omega}-2AB\sigma^{\omega}_{H}]\right\}\\ +\frac{i}{3\tau_{H}}\left[\left(\frac{2\omega}{T^{2}}\right)^{2}\dfrac{6}{T^{2}}\left(\sigma_{A}^{\omega}-\frac{T^{2}}{3}\Delta H\right)[(\sigma_{A}^{\omega})^{2}-(\sigma_{H}^{\omega})^{2}]\right.\\ -\left.\frac{\kappa_{\Omega}^{2}A}{H}\dfrac{6}{T^{2}}(A\sigma_{A}^{\omega}-B\sigma_{H}^{\omega})\right]=0\,.$ (164)
We can solve the previous equation perturbatively in $\kappa_{\Omega}$, considering the following expansion for $\omega$:
@@
$\left(\frac{2\omega_{0}}{T^{2}}\right)^{2}\left(\sigma^{\omega}_{A}-\frac{T^{2}}{3}\Delta H\right)[(\sigma^{\omega}_{A})^{2}-(\sigma^{\omega}_{H})^{2}]\\ \times\left(\frac{2\omega_{0}}{T^{2}}+\dfrac{2i}{\tau_{H}T^{2}}\right)=0\,.$ (166)
There is a non-trivial and purely dissipating solution $\omega_{0}=-i/\tau_{H}$ and a double trivial root $\omega_{0}=0$. We now compute the corrections to the trivial root from the next-to-leading order for the master equation:
$\dfrac{2i}{\tau_{H}T^{2}}\left[\left(\frac{2\omega_{1}}{T^{2}}\right)^{2}\left(\sigma_{A}^{\omega}-\frac{T^{2}}{3}\Delta H\right)[(\sigma_{A}^{\omega})^{2}-(\sigma_{H}^{\omega})^{2}]\right.$
$\left.-\frac{A}{H}(A\sigma_{A}^{\omega}-B\sigma_{H}^{\omega})\right]\kappa_{\Omega}^{2}=0$ (167)
@@
$\omega_{1}^{\pm}=\pm\dfrac{T^{2}}{2}\sqrt{\dfrac{A(A\sigma_{A}^{\omega}-B\sigma_{H}^{\omega})}{H\left(\sigma_{A}^{\omega}-\frac{T^{2}}{3}\Delta H\right)[(\sigma_{A}^{\omega})^{2}-(\sigma_{H}^{\omega})^{2}]}}\,.$ (168)
The quantity $A(A\sigma_{A}^{\omega}-B\sigma_{H}^{\omega})$ is negative semi-definite, whereas the remaining factor inside the square root is positive semi-definite. Specifically, from the definitions in Sec. 2.4, we find $\sigma_{A}^{\omega}>\sigma_{H}^{\omega}$ and
$H=\frac{15\alpha_{V}^{4}+30\pi^{2}\alpha_{V}^{2}+7\pi^{4}}{\pi^{2}(15\alpha_{V}^{2}+7\pi^{2})}$ $>0,$
$\frac{1}{T^{2}}\left(\sigma_{A}^{\omega}-\frac{T^{2}}{3}\Delta H\right)=\frac{15\alpha_{V}^{4}+6\alpha_{V}^{2}\pi^{2}+7\pi^{4}}{6\pi^{2}(15\alpha_{V}^{2}+7\pi^{2})}$ $>0,$ (169)
@@
$\dfrac{1}{T^{2}}A(A\sigma_{A}^{\omega}-B\sigma_{H}^{\omega})\simeq-0.000432\alpha_{V}^{2}<0.$ (170)
@@
Figure 13: Explicit demonstration that the left-hand side of Eq. (170) is negative and therefore there is always one unstable mode at small wavenumber when charge dissipation is implemented on the basis of charge fluctuations (see Appendix A).
@@ Appendix B Helicity relaxation time at finite chemical potential
In this section of the appendix, we extend the calculation of the helicity relaxation time $\tau_{H}$ presented in Ref. Ambrus:2019khr to the case of an unpolarized, charged plasma, i.e. at finite vector chemical potential $\mu_{V}=\alpha_{V}T$. We start from Eqs. (101), (102) and (104) of Ref. Ambrus:2019khr :
@@
with $\tilde{f}^{\sigma}_{{\mathbf{p}},\lambda}=1-f^{\sigma}_{{\mathbf{p}},\lambda}$ and $s=E^{2}_{\rm cm}=(p+k)^{2}=(p^{\prime}+k^{\prime})^{2}$ being the center-of-mass energy, while $\theta_{\rm cm}$ is the angle between ${\mathbf{p}}^{\prime}$ and ${\mathbf{p}}$, measured in the center of mass frame, satisfying
@@
where $f^{\sigma}_{0{\mathbf{p}}}=[e^{\beta E_{\mathbf{p}}-\sigma\alpha_{V}}+1]^{-1}$ is the equilibrium distribution for a charged, unpolarized fluid, thus extending Eq. (105) of Ref. Ambrus:2019khr for the neutral fluid to the case of a charged fluid. Note that in Eq. (107) of Ref. Ambrus:2019khr , Eq. (171) was replaced by $dQ_{H}/dt=-Q_{H}/\tau_{H}$, which is valid under the assumption of a neutral background with a small helicity imbalance, i.e. $\mu_{V}=\mu_{A}=0$ and $\lvert\mu_{H}\rvert\ll T$, when $Q_{H}\simeq\mu_{H}T^{2}/3$. In this work, we consider the case when $\mu_{V}$ is arbitrary, such that instead Eq. (171) becomes
@@
where we used the property $\tilde{f}^{\sigma}_{0{\mathbf{p}}}=1-f^{\sigma}_{0{\mathbf{p}}}$. Writing now as in Eq. (159) of Ref. Ambrus:2019khr ,
@@
Figure 14: Ratio $\tau_{H}(0)/\tau_{H}(\alpha_{V})$ between the helicity relaxation time in a neutral plasma and at dimensionless vector chemical potential $\alpha_{V}=\mu_{V}/T$, in an unpolarized plasma. The inset plot shows the agreement between the numerical integration and the fitted function in a log-log scale.
The integral with respect to $\varphi_{{\mathbf{p}}^{\prime}}$ can be performed using Eq. (172) to replace $1-\cos\theta_{\rm cm}$. After boosting $p^{\prime}$ to the center of mass frame, $p^{\mu}$ is replaced by $p^{\mu}_{\rm cm}=(E_{\rm cm}/2,{\mathbf{p}}_{\rm cm})$, with
@@
Splitting ${\mathbf{p}}_{\rm cm}={\mathbf{p}}_{\lvert\rvert}+{\mathbf{p}}_{\perp}$, where ${\mathbf{p}}_{|\rvert}$ is oriented along ${\mathbf{k}}+{\mathbf{p}}$ and ${\mathbf{p}}_{\perp}$ is perpendicular to this vector, we have
$p_{\lvert\rvert}=\frac{{\mathbf{p}}_{\rm cm}\cdot({\mathbf{p}}+{\mathbf{k}})}{\lvert{\mathbf{p}}+{\mathbf{k}}\rvert}=\frac{E_{\rm cm}(p^{0}-k^{0})}{2\lvert{\mathbf{p}}+{\mathbf{k}}\rvert},$ (182)
@@
$\tau_{H}$ $=0.392\times\frac{\pi^{3}\beta}{N_{f}\alpha_{\rm QCD}^{2}}$
$\simeq\left(\frac{250\ {\rm MeV}}{k_{B}T}\right)\left(\frac{1}{\alpha_{\rm QCD}}\right)^{2}\left(\frac{2}{N_{f}}\right)\times 4.80\ {\rm fm}/c.$ (185)
@@
$\tau_{H}^{-1}(\alpha_{V}\gg 1)=\tau^{-1}_{\rm deg.}\frac{e^{-\lvert\alpha_{V}\rvert}}{\lvert\alpha_{V}\rvert},$ (186)
where the best fit was found to be given by $\tau^{-1}_{\rm deg.}=118.6\ c/{\rm fm}$, or $\tau_{\rm deg.}\simeq 8.432\times 10^{-3}\,{\rm fm}/c$.